David’s DVD Repair Company wants to open a repair shop in a suburb of a major...

  1. David’s DVD Repair Company wants to open a repair shop in a suburb of a major metropolitan area. The industry association estimates that 30% of DVD players are repaired by similar service companies and that the average owner spends $80 per DVD player on repairs each year. The available commerce data indicates that there are 50,000 DVD players in the metropolitan service area. Three other competitors exist within a 25 mile radius of the proposed business location. Based on a consumer survey, the owner believes that he can capture 20% of the market in the first year of operation.

Based on this data, address the following requirements:

What is the potential number of DVD players likely to be commercially repaired during one year?

How much revenue can David’s DVD Repair Company expect to generate in its first year of operations?
Answer #1

1) Potential number of DVD players likely to be commercially repaired during one year = 50,000 * 30% = 15,000.

2) Amount of revenue David's DVD Repair Company expects to generate in its first year of operations = 15,000 * 20% * $80 = $2,40,000.